Solve for the new pressure in each of the following, with n and V constant:A gas with a pressure of 1.05 atm at 90 ∘C is cooled to -18 ∘C A sample of N2 with a pressure of 790 mmHg at -85 ∘C is heated to 33 ∘C.
Answer #1

1)

Given:

Pi = 1.05 atm

Ti = 90.0 oC

= (90.0+273) K

= 363 K

Tf = -18.0 oC

= (-18.0+273) K

= 255 K

use:

Pi/Ti = Pf/Tf

1.05 atm / 363.0 K = Pf / 255.0 K

Pf = 0.7376 atm

Answer: 0.738 atm

B)

Given:

Pi = 790 mmHg

Ti = -85.0 oC

= (-85.0+273) K

= 188 K

Tf = 33.0 oC

= (33.0+273) K

= 306 K

use:

Pi/Ti = Pf/Tf

790 mmHg / 188.0 K = Pf / 306.0 K

Pf = 1286 mmHg

= 1286/760 atm

= 1.69 atm

Answer: 1.69 atm